# A Unified Multi-Wavelength Data Analysis Workflow with gammapy

### v1

@@ Abstract
Context. The Flat Spectrum Radio Quasar OP 313 entered an enhanced activity phase in November 2023 and has undergone multiple flares since then which have motivated the organisation of several large multi-wavelength campaigns, including two deep observations from the hard X-ray telescope NuSTAR. The broadband emission from OP 313 during these two observations is investigated under a new unified analysis framework, with data spanning from optical to $\gamma$ rays.
Aims. Traditional methods for analyzing blazar emissions often rely on proprietary software tailored to specific instruments, making it challenging to integrate and interpret data from multi-wavelength campaigns comprehensively. This study demonstrates the feasibility of utilizing gammapy, an open-source Python package, and common data formats originally developed for $\gamma$-ray instrumentation, to perform a consistent multi-instrument analysis. This enables a forward folding approach that fully incorporates source observations, detector responses, and various instrumental and astrophysical backgrounds. The methodology is applied as an example to recent data collected from the distant quasar OP 313.
@@ 1 Introduction
Active Galactic Nuclei (AGNs), in particular those with large-scale jets of ultra-relativistic particles, are the predominant extragalactic sources of $\gamma$-ray photons. Their non-thermal emission extends over the entire electromagnetic spectrum, from radio to $\gamma$ rays, yet the driving physical mechanism that explains the radiation in each band is different. At low energies, from radio to optical (in some cases up to X-rays), the emission is often assumed to be mainly due to synchrotron radiation from relativistic electrons (Konigl 1981). At higher energies, the situation is less clear and diverse mechanisms are assumed to explain the emission, from inverse Compton of the same electrons with the synchrotron radiation (synchrotron-self-Compton or SSC, see Maraschi et al. 1992) or with external thermal photons from the AGN structure (external Compton or EC, e.g. Dermer & Schlickeiser 1993), to various hadronic processes: proton-synchrotron, proton-proton (see e.g. Aharonian 2002).
@@
However, detecting very-high-energy (VHE; E $>$ 100 GeV) $\gamma$-ray emission from AGNs — especially for those found at high redshifts — is challenging due to attenuation by the extragalactic background light (EBL, e.g. Domínguez et al. 2011; Saldana-Lopez et al. 2021), which absorbs $\gamma$-rays from distant sources and is an irreducible background that encodes important information about the star formation history in the Universe. The induced absorption produces an energy-dependent imprint on the blazar spectrum, reducing significantly the observable $\gamma$-ray flux in the VHE band. At the same time, it provides an opportunity to constrain the density of the EBL indirectly, provided that we can infer the intrinsic spectrum of the source. For sources at redshift $z\sim 1$ and sub-TeV photons, the most relevant part of the spectrum of the EBL is the so-called Cosmic Optical Background component.
@@
In this work, we aim to validate a new analysis and data management workflow based on the standard format proposed by the Data formats for $\gamma$-ray astronomy (Nigro et al. 2021) and the open-source tool gammapy (Donath et al. 2023), and provide a working prototype data archive for the source, retaining full instrument response metadata for reproducibility in future analyses and interpretation of these observations. This manuscript focuses on demonstrating and validating a new methodology for multi-wavelength analysis. For this demonstration we combined datasets from a small coordinated multi-wavelength follow-up campaign on OP 313 developed on March 4th 2024 (MJD60373) and March 15th (MJD60384), utilizing public data from Swift-UVOT, Swift-XRT, NuSTAR, and Fermi-LAT, and data from Liverpool IO:O in the optical regime. The resulting data products, in standard OGIP/GADF formats compatible with gammapy, are readily available in GitHub [^1] https://github.com/mireianievas/gammapy_mwl_workflow and described in more details in Appendix A. We leave the integration of this dataset with a larger campaign, including proprietary VHE $\gamma$-ray data from LST and MAGIC, as well as additional optical, infrared, and radio results, for a separate work coordinated by the CTAO LST Collaboration. To keep the main body of the manuscript as focussed as possible, we will refer always to data and results from the first observing night on March 4th, 2024 (MJD60373), and comment briefly on the results of the application of the analysis framework to the second night on March 15th, 2024 (MJD60384) in Appendix E.
@@
Our proposed method involves a forward modeling approach within a unified framework for all multi-wavelength datasets, covering optical (z band) to HE $\gamma$-rays, spanning nearly twelve orders of magnitude in energy. The mathematical representation of this method is as follows:
@@
$\begin{split}N(E,X)=\sum_{sources}\mathcal{R}&(E,E_{\rm true},X)\ast\mathcal{P}(E_{\rm true},X,X_{true})\\ \ast&\left[\mathcal{E}(E_{\rm true},X_{true})\times\phi(E_{\rm true})\right],\end{split}$ (1)
@@
where the number of expected excess event density $N(E,X)$ (in space and energy) is given as a function of the sum over a number of finite emitters of the differential spectrum $\phi(E_{\rm true})$ of each source times the exposure at its location, convoluted with the PSF $\mathcal{P}(E_{\rm true},X,X_{true})$, which informs about the actual distribution of the measured counts given a location in the detector/sky, possibly as a function of energy, and a redistribution or migration matrix $\mathcal{R}(E,E_{\rm true},X)$ which provides the relation between the theoretical energy and the measured energy of the event, as a function of location in the detector/sky.
@@
The manuscript is structured as follows: Section 2 details the motivation behind developing a new method for analyzing and archiving high-level data across multiple bands. Section 3 describes the different instruments and procedures used to build the datasets and Instrument Response Functions (IRFs). Section 4 presents the validation of the data formats and analysis methods, along with a phenomenological model of the emission of the source. The discussion of limitations, possible extensions, and future work is included in Section 5, and the main results are summarized in Section 6. Wherever applicable, we use a flat $\mathrm{\Lambda}$CDM cosmology, with $\mathrm{H_{0}=70\,km\,s^{-1}\,Mpc^{-1}}$, $\mathrm{\Omega_{M}}=0.3$, and $\mathrm{T_{CMB,0}=2.725\ K}$.
@@ 2 Motivation
Firstly, storing flux points in tables or plots often results in significant information loss. This approach typically fails to account for the energy resolution of the instruments if no unfolding (Schmitt 2017) is performed, a process that is inherently ill-posed. Consequently, the flux points may become correlated. Additionally, errors are commonly treated as Gaussian, which is inappropriate for instruments operating in low-count regimes where Poissonian statistics are more suitable (see e.g. Yamada et al. 2019). ULs are frequently represented as single values at specific confidence levels, resulting in the loss of information contained in the underlying probability distributions. This practice can lead to biased interpretations, especially when statistical fluctuations, ULs and non-detections are ignored (e.g. Essey & Kusenko 2014).
@@
Moreover, current methods fall short in handling multiplicative models in SED flux points, such as correcting for hydrogen column density ($N_{H}$) in X-ray data, extinction in optical/UV or EBL absorption in the $\gamma$-ray regime. Typically, these corrections are applied to the reconstructed flux points, which can hinder accurate statistical correction for absorption components in the SED, particularly if the correction is applied differently for each dataset (this is sometimes the case for optical instruments).
@@
Recent advancements have explored moving away from flux points towards forward-folding techniques, which have demonstrated promising results in characterizing the EBL (Acciari et al. 2019) using VHE data. This forward-folding approach offers a more comprehensive statistical analysis of multi-wavelength datasets, addressing many shortcomings of traditional flux-point methods. Although forward-folding is inherently model-dependent, it provides greater stability compared to unfolding methods while solving the same underlying problem. By incorporating forward modeling, emission and absorption models can be seamlessly integrated into the analysis, enabling better handling of uncertainties in parameters such as hydrogen column density and EBL. In this work we address one of the main drawbacks of the forward-folding method, which is how to distribute the actual data together with a consistent description of the instrument that allows to interpret it without the need to install the software for each instrument.
@@
Our immediate goal is therefore to integrate high-level data from various photon-counting experiments into a unified analysis framework. To facilitate this, we propose using the standardized format for high-level data products, which ensures easier distribution of the data and instrument description, as well as reproducibility of the results. gammapy, an open-source Python package based on the popular libraries numpy (Harris et al. 2020) and astropy (Astropy Collaboration et al. 2013, 2018, 2022), is particularly well-suited for this task. As the official science tool for the Cherenkov Telescope Array Observatory (CTAO), gammapy is actively developed and compatible with both OGIP (X-ray) and GADF ($\gamma$-ray) data formats. This makes it an ideal choice for our analysis, enabling the joint analysis of different types of datasets, whether one-dimensional (1D) if only a distribution of counts as a function of energy is stored, or three-dimensional (3D) where data cubes of three dimensions (two spatial coordinates and one energy coordinate) are available. Finally, we show that even in the case of single-channel photometric datasets the proposed format is well suited to describe the data.
@@
Furthermore, gammapy allows the incorporation of physically motivated models through external emission libraries — including synchrotron, inverse Compton, absorption components such as EBL absorption — or even extend it with models from the X-ray library sherpa (Freeman et al. 2001; Doe et al. 2007) to add hydrogen absorption, and interstellar extinction. Public radiation codes like agnpy (Nigro et al. 2022), jetset (Tramacere et al. 2009, 2011; Tramacere 2020), and naima (Zabalza 2015) offer a more accurate representation of underlying physical processes compared to traditional functional or empirical models available in standard X-ray analysis packages like xspec (Arnaud 1996) and sherpa (Freeman et al. 2001; Doe et al. 2007). This capability allows for a more accurate and statistically robust comparison of different models.
@@ 3.1.1 Data reduction
We collected Fermi-LAT Pass 8 SOURCE class events with energies larger than 100100100\,100MeV within a region of interest (ROI) of 151515\,15deg radius around the OP 313 position from October 1, 2023, to April 1, 2024 (183 days). Night-by-night binned analyses (centered at 00:00 UTC were performed with the Fermi Science Tools (Fermi Science Support Development Team 2019), using enrico (Sanchez & Deil 2015) as a high-level wrapper to manage the individual jobs. We utilized the latest available version of LAT’s IRFs (P8R3_SOURCE_V3) and applied a conservative zenith cut of 909090\,90deg to avoid Earth’s limb contamination and a DATA_QUAL==1 && LAT_CONFIG==1 filter to ensure that only good quality events were analyzed, following the recommendations in Cicerone [^2] https://fermi.gsfc.nasa.gov/ssc/data/analysis/documentation/Cicerone/Cicerone_Likelihood/Exposure.html.
The sky model for the analysis was generated using the Fermi-LAT 14-Year Point Source Catalog (4FGL-DR4 Abdollahi et al. 2022; Ballet et al. 2023), incorporating sources within a 202020\,20deg radius around OP 313. Spectral parameters for weak sources ($<10\,\sigma$ significance) and distant sources ($>5\,$deg) were fixed, except for normalizations. For OP 313, and because we only integrate 242424\,24h per analysis, we adopted a simple power-law (PWL) spectral model, including the EBL absorption model from Domínguez et al. (2011). Positions and extensions of all sources from the 4FGL-DR4 were fixed, leaving only the spectral parameters of OP 313 (normalization and spectral index) and the normalization of the isotropic background component as free parameters.
@@ 3.1.2 Long-term analysis of the high-energy flare
In addition to focusing on two specific nights with comprehensive multi-wavelength coverage, we examined the broader activity of OP 313 to contextualize its emission during these nights within a longer-term flare that has persisted for several months in 2023-2024. Figure 1 shows the evolution of the nightly flux above 100100100\,100MeV for OP 313, with 24-hour bins centered at midnight UTC. It also depicts the evolution of the spectral index in the Fermi-LAT band and its correlation with the estimated flux, hinting that during high emission states, the spectrum of OP 313 becomes harder, with an index in the LAT band smaller than $\sim 2$. From Figure 1, we observe that the source entered an intermediate $\gamma$-ray activity plateau phase following a very strong flare at the end of February.
@@ 3.1.3 Dataset and IRFs
    Initial Model Map convolved with the IRFs to produce the predicted counts (npred map) and again integrated over the entire energy range just for visualization purposes. The original, 3D version, of this Model is fit to the data (‘counts cube’) by gammapy.
@@
    Containment Radius: 1D representation of the PSF corresponding to $68\%$ and $95\%$ of the events contained as a function of energy, indicating the spatial resolution of the LAT. As opposed to the native analysis, Fermitools, our conversion to gammapy format assumes a non-varying PSF. For a bright point-like source analysis in a 202020\,20deg region size, we estimate the impact of not modelling the PSF correctly at the edges to be negligible.
@@
    Energy Dispersion: Shows the distribution of reconstructed energies versus true energies, providing information on how accurately the instrument measures the energy of $\gamma$-ray events and possible energy biases (mostly present at the lowest energies, $E\lesssim 100\,$MeV).
@@
    Exposure at the FoV Center: 1D projection of the exposure as a function of true energy at the center of the field of view, showing the energy dependence of the LAT’s sensitivity.
Figure 2: Fermi-LAT dataset description for the first selected night. From top left to bottom right: (a) Integral counts map, smoothed with a Gaussian kernel of $0.2$ deg. (b) Initial model map, convolved with the IRFs (npred map). (c) $68\%$ and $95\%$ containment radius as a function of energy. (d) Energy dispersion (reconstructed energy as a function of true energy). (e) Exposure map at 191.5191.5191.5\,191.5GeV. (f) Exposure at the FoV center as a function of true energy.

### v2

@@ Abstract
Context. The Flat Spectrum Radio Quasar OP 313 entered an enhanced activity phase in November 2023 and has undergone multiple flares since then, which have motivated the organization of several large multi-wavelength campaigns, including two deep observations from the hard X-ray telescope NuSTAR. The broadband emission from OP 313 during these two observations is investigated under a new unified analysis framework, with data spanning from optical to $\gamma$ rays.
Aims. Traditional methods for analyzing blazar emission often rely on proprietary software tailored to specific instruments, making it challenging to integrate and interpret data from multi-wavelength campaigns comprehensively. This study demonstrates the feasibility of utilizing gammapy, an open-source Python package, and common data formats originally developed for $\gamma$-ray instrumentation, to perform a consistent multi-instrument analysis. This enables a forward folding approach that fully incorporates source observations, detector responses, and various instrumental and astrophysical backgrounds. The methodology is applied as an example to recent data collected from the distant quasar OP 313.
@@ 1 Introduction
Active Galactic Nuclei (AGNs), in particular those with large-scale jets of ultra-relativistic particles, are the predominant extragalactic sources of $\gamma$-ray photons. Their non-thermal emission extends over the entire electromagnetic spectrum, from radio to $\gamma$ rays; yet the driving physical mechanism that explains the radiation in each band is different. At low energies, from radio to optical (in some cases up to X-rays), the emission is often attributed to synchrotron radiation from relativistic electrons (Konigl 1981). At higher energies, the situation is less clear and diverse mechanisms are assumed to explain the emission, from inverse Compton of the same electrons with the synchrotron radiation (synchrotron-self-Compton or SSC, see Maraschi et al. 1992) or with external thermal photons from the AGN structure (external Compton or EC, e.g. Dermer & Schlickeiser 1993), to various hadronic processes: proton-synchrotron, proton-proton (see e.g. Aharonian 2002).
@@
However, detecting very-high-energy (VHE; E $>$ 100 GeV) $\gamma$-ray emission from AGNs — especially for those found at high redshifts — is challenging due to attenuation by the extragalactic background light (EBL, e.g. Domínguez et al. 2011; Saldana-Lopez et al. 2021), which absorbs $\gamma$ rays from distant sources and is an irreducible background that encodes important information about the star formation history in the Universe. The induced absorption produces an energy-dependent imprint on the blazar spectrum, reducing significantly the observable $\gamma$-ray flux in the VHE band. At the same time, it provides an opportunity to constrain the density of the EBL indirectly, provided that we can infer the intrinsic spectrum of the source. For sources at redshift $z\sim 1$ and sub-TeV photons, the most relevant part of the spectrum of the EBL is the so-called Cosmic Optical Background component.
@@
In this work, we aim to validate a new analysis and data management workflow using the OGIP format [^1] https://heasarc.gsfc.nasa.gov/docs/heasarc/ofwg/docs/spectra/ogip_92_007.pdf (Corcoran et al. 1995), the Gamma Astro Data Formats (GADF) initiative (Deil et al. 2017; Nigro et al. 2021), and the open-source tool gammapy (Donath et al. 2023), and provide a working prototype data archive for the source, retaining instrument response metadata for reproducibility in future analyses and interpretation of these observations. This manuscript focuses on demonstrating and validating a novel methodology for multi-wavelength analysis. For this demonstration we combined datasets from a small coordinated multi-wavelength follow-up campaign on OP 313 developed on March 4, 2024 (MJD60373) and March 15, 2024 (MJD60384), utilizing public data from Swift-UVOT, Swift-XRT, NuSTAR, and Fermi-LAT, and data from Liverpool IO:O in the optical regime. The resulting data products, in standard OGIP/GADF formats compatible with gammapy, are readily available in Zenodo (Nievas Rosillo et al. 2024) and GitHub [^2] https://github.com/mireianievas/gammapy_mwl_workflow and described in more detail in Appendix A. We defer the integration of this dataset with a larger campaign, including proprietary VHE $\gamma$-ray data from LST-1 and MAGIC, as well as additional optical, infrared, and radio results, for a separate work coordinated by the CTAO LST Collaboration. To keep the main body of the manuscript as focused as possible, we will refer always to data and results from the first observing night on March 4, 2024 (MJD60373), and comment briefly on the results of the application of the analysis framework to the second night on March 15, 2024 (MJD60384) in Appendix F.
@@
Our proposed method involves a forward modeling approach within a unified framework for all multi-wavelength datasets, covering near-infrared (z-band) to high-energy (HE, $100\,\mathrm{MeV}<E<100\,\mathrm{GeV}$) $\gamma$-rays, spanning nearly 11 orders of magnitude in energy. The mathematical representation of this method is as follows:
@@
$\begin{split}N(E,X)=\sum_{{\rm sources}}\mathcal{R}&(E,E_{\rm true},X)\ast\mathcal{P}(E_{\rm true},X,X_{\rm true})\\ \ast&\left[\mathcal{E}(E_{\rm true},X_{\rm true})\times\phi(E_{\rm true})\right],\end{split}$ (1)
@@
where the number of expected excess event density $N(E,X)$ (in space and energy) is given as a function of the sum over a number of finite emitters of the differential spectrum $\phi(E_{\rm true})$ of each source times the exposure $\mathcal{E}(E_{\rm true},X_{\rm true})$ at its location, convoluted with the PSF $\mathcal{P}(E_{\rm true},X,X_{\rm true})$, which informs about the actual distribution of the measured counts given a location in the detector/sky, possibly as a function of energy, and a redistribution or migration matrix $\mathcal{R}(E,E_{\rm true},X)$ which provides the redistribution between the true energy and the measured energy of the event, as a function of the location in the detector $X$ and the sky $X_{\rm true}$.
@@
The manuscript is structured as follows: Section 2 details the motivation for developing a new method for analyzing and archiving high-level data across multiple bands. Section 3 describes the different instruments and procedures used to build the datasets and Instrument Response Functions (IRFs). Section 4 presents the validation of the data formats and analysis methods, along with a phenomenological model of the emission of the source. The discussion of limitations, possible extensions, and future work is included in Section 5, and the main results are summarized in Section 6. Wherever applicable, we use a flat $\mathrm{\Lambda}$CDM cosmology, with $\mathrm{H_{0}=70\,km\,s^{-1}\,Mpc^{-1}}$, $\mathrm{\Omega_{M}}=0.3$, and $\mathrm{T_{CMB,0}=2.725\ K}$.
@@ 2 Motivation
First, storing flux points in tables or plots often results in significant information loss. This approach typically fails to account for the energy resolution of the instruments if no unfolding (Schmitt 2017) is performed, a process that is inherently ill-posed. Consequently, the flux points may become correlated. Additionally, errors are commonly treated as Gaussian, which is inappropriate for instruments that operate in low-count regimes where Poisson statistics are more suitable (see e.g. Yamada et al. 2019). ULs are frequently represented as single values at specific confidence levels, resulting in the loss of information contained in the underlying probability distributions. This practice can lead to biased interpretations, especially when statistical fluctuations, ULs and non-detections are ignored (e.g. Essey & Kusenko 2014).
@@
Moreover, current methods fall short in handling multiplicative models in SED flux points, such as correcting for hydrogen column density ($N_{H}$) in X-ray data, Galactic extinction in optical/UV or EBL absorption in the $\gamma$-ray regime. Typically, these corrections are applied to the reconstructed flux points, which can hinder accurate statistical correction for absorption components in the SED, particularly if the correction is applied differently for each dataset (this is sometimes the case for works based on archival data samples, e.g. Nievas Rosillo et al. 2022).
@@
Recent advancements have explored moving away from flux points towards forward-folding techniques, which have demonstrated promising results in characterizing the EBL (Acciari et al. 2019) using VHE data. This forward-folding approach offers a more comprehensive statistical analysis of multi-wavelength datasets, addressing many shortcomings of traditional flux-point methods. One key distinction is that flux-point fitting loses access to instrument-specific information (such as energy resolution, effective area, instrumental backgrounds, and PSF) while forward-folding incorporates these details as an essential part of the analysis convolving complete emission models with the IRFs to calculate the likelihood. By adopting the forward modeling technique, emission and absorption models can be seamlessly integrated into the analysis, enabling better handling of uncertainties in parameters such as hydrogen column density and EBL. Despite its advantages, forward-folding presents practical challenges, such as the need to install and configure analysis tools for each instrument and manage the orchestration of communication between them to compute the multi-instrument likelihood.
@@
Our immediate goal is therefore to integrate high-level data from various photon-counting experiments into a unified analysis framework. To facilitate this, we propose using the standardized format for high-level data products, which ensures easier distribution of the data and instrument description, as well as reproducibility of the results. gammapy (Acero et al. 2024), an open-source Python package based on the popular libraries numpy (Harris et al. 2020) and astropy (Astropy Collaboration et al. 2013, 2018, 2022, 2024), is particularly well-suited for this task. As the official science tool for the Cherenkov Telescope Array Observatory (CTAO), gammapy is actively developed and compatible with both OGIP (X-ray) and GADF ($\gamma$-ray) data formats. This makes it an ideal choice for our analysis, enabling the joint analysis of different types of datasets, whether one-dimensional (1D) if only a distribution of counts as a function of energy is stored, or three-dimensional (3D) where data cubes of three dimensions (two spatial coordinates and one energy coordinate) are available. Finally, we show that even in the case of single-channel photometric datasets the proposed format is well suited to describe the data.
@@
Furthermore, gammapy allows the incorporation of physically motivated models through external emission libraries — including synchrotron, inverse Compton, absorption components such as EBL absorption — or even extend it with models from the X-ray library sherpa (Freeman et al. 2001; Doe et al. 2007; Burke et al. 2024) to add hydrogen absorption, and interstellar extinction. Public radiation processes modeling codes like agnpy (Nigro et al. 2022, 2023), jetset (Tramacere et al. 2009, 2011; Tramacere 2020), and naima (Zabalza 2015) offer a more accurate representation of underlying physical processes compared to traditional functional or empirical models available in standard X-ray analysis packages like xspec (Arnaud 1996) and sherpa. This capability allows for a more accurate and statistically robust comparison of different models.
@@ 3.1.1 Data reduction
We collected Fermi-LAT Pass 8 SOURCE class events with energies larger than 100100100\,100MeV within a region of interest (ROI) of 151515\,15deg radius around the OP 313 position from October 1, 2023, to April 1, 2024 (183 days). Nightly binned analyses, centered at 00:00 UTC, were performed with the Fermi Science Tools (Fermi Science Support Development Team 2019), using enrico (Sanchez & Deil 2015) as a high-level wrapper to manage the individual jobs. We utilized the latest available version of LAT’s IRFs (P8R3_SOURCE_V3) and applied a conservative zenith cut of 909090\,90deg to avoid Earth’s limb contamination and a DATA_QUAL==1 && LAT_CONFIG==1 filter to ensure that only good quality events were analyzed, following the recommendations in Cicerone [^3] https://fermi.gsfc.nasa.gov/ssc/data/analysis/documentation/Cicerone/.
The sky model for the analysis was generated using the Fermi-LAT 14-Year Point Source Catalog (4FGL-DR4 Abdollahi et al. 2022; Ballet et al. 2023), incorporating sources within a 202020\,20deg radius around OP 313. Because we only integrate 242424\,24h per analysis, we adopted a simple power-law (PWL) spectral model for OP 313, including the EBL absorption model from Domínguez et al. (2011). Positions and extensions of all sources from the 4FGL-DR4 were kept fixed, leaving only the spectral parameters of OP 313 (normalization and spectral index) and the normalization of the isotropic background component as free parameters. Spectral parameters for weak sources ($<10\,\sigma$ significance) and distant sources ($>5\,$deg) were kept fixed, except for normalizations.
@@ 3.1.2 Long-term analysis of the high-energy flare
In addition to focusing on two specific nights with comprehensive multi-wavelength coverage, we examined the broader activity of OP 313 to contextualize its emission during these nights within a longer-term flare that persisted over several months in 2023-2024. Figure 1 shows the evolution of the nightly flux above 100100100\,100MeV for OP 313, with 24-hour bins centered at midnight UTC. It also depicts the evolution of the spectral index in the Fermi-LAT energy band and its correlation with the estimated flux, hinting that during high emission states, the spectrum of OP 313 becomes harder, with an index in the LAT band smaller than $\sim 2$. Figure 1 shows that the source entered an intermediate $\gamma$-ray activity plateau phase following a very strong flare at the end of February.
@@ 3.1.3 Dataset and IRFs
    Initial Model Map convolved with the IRFs to produce the predicted counts (npred map) and again integrated over the entire energy range just for visualization purposes. The original 3D version of this Model is fit to the data (‘counts cube’) by gammapy.
@@
    Containment Radius: 1D representation of the PSF corresponding to $68\%$ and $95\%$ of the events contained as a function of energy, indicating the spatial resolution of the LAT. As opposed to the native analysis, Fermitools, our conversion to gammapy format assumes a non-varying PSF. For the analysis of a bright point-like source in a $20^{\circ}$ region, the impact of spatial variations of the PSF of LAT is negligible.
@@
    Energy Dispersion: Distribution of reconstructed energies versus true energies, providing information on how accurately the instrument measures the energy of $\gamma$-ray events and energy bias (significant at the lowest energies, $E\lesssim 100\,$MeV).
@@
    Exposure at the FoV center: 1D projection of the exposure as a function of true energy at the center of the field of view, showing the energy dependence of LAT’s sensitivity.
Figure 2: Fermi-LAT dataset description for the first selected night. From top left to bottom right: (a) Integral counts map, smoothed with a Gaussian kernel of $0.2^{\circ}$. (b) Initial model map, convolved with the IRFs (npred map). (c) $68\%$ and $95\%$ containment radius as a function of energy. (d) Energy dispersion (reconstructed energy as a function of true energy). (e) Exposure map at 10.610.610.6\,10.6GeV. (f) Exposure at the FoV center as a function of true energy.